# Evaluating the Limits of QAOA Parameter Transfer at High-Rounds on Sparse Ising Models With Geometrically Local Cubic Terms

### v1

@@ III.1 Classical QAOA Angle Learning with Statevector Simulations
Figure 1 plots the numerical quantities of the learned QAOA angles $\beta,\gamma$ for the ensemble of $12$ different $16$-qubit heavy-hex spin glass instances, learned until $p=49$ or the mean (noiseless) energy is at least $0.01$ (energy) below the optimal ground-state energy. During training, the QAOA angles are not restricted or bounded to a certain range. Figure 2 plots the same but for the ensemble of $27$ qubit instances. This set of $24$ QAOA angles are the parameters that are used in the entirety of this study for evaluating how well, or not, parameter transfer works up to these very high depth QAOA circuits.
Figure 4: Numerical simulation of parameter transfer of one of the $16$ qubit problem instances onto the other $11$ $16$-qubit problem instances (left) and onto the $12$ $27$-qubit problem instances (right). These simulations are all noiseless (no shot noise). y-axis is the approximation ratio for the given target problem instances, x-axis is $p$ QAOA steps. Each line corresponds to a different problem instance (which were not using in the training of this set of fixed angles), which is indexed in the legend numerically and by the all positive and negative coefficient Ising models. Red star markers denote the $p$ step where the expectation value decreases in quality (increases in energy) for that given problem instance, if that occurs. 88 out the $11$ instances in the left plot have at least one non-improving $p$ step, and 99 out of $12$ of the instances in the right plot have at least one non-improving $p$ step. However, there is an overall trend for all problem instances shown in these plots that the higher $p$ angles do give consistently improving energies, even if there are some blips of non-monotonic improvements.
Our trained angles all lie in the range $0,2\pi)$, however, there are several periodicities and symmetries that enable us to further transfer the angles into mathematical equivalent angles plotted in Figures 12:
@@
  • Lastly, some of the trained QAOA angles have very different behavior that is not seen in the other problem instances – namely, instance number 5 in Figure 1 shows highly fluctuating angles for the largest (converged) $p$ that was found. This highly fluctuating instance could be related to the convergence to nearly an approximation ratio of $1$.
@@
More generally, these optimized QAOA schedules closely follow recent studies that have drawn connections between highly optimized QAOA trajectories, and continuous-time quantum annealing schedules, which typically result in smooth QAOA angle trajectories [82, 83, 48, 84, 85] (this connection can be understood as QAOA being a type of discretized of continuous-time adiabatic quantum annealing). These highly optimized QAOA angles, shown in Figures 1 and 2, therefore provide more evidence for this fundamental connection between continuous time adiabatic evolution and the discretized QAOA algorithm.
@@
Figure 3 plots the mean noiseless approximation ratios for the set of trained QAOA angles shown in Figures 1 and 2, evaluated on the problem instances on which the QAOA parameters were trained. Figure 3 shows the rate of energy improvement as a function of $p$ (which quickly plateaus at large $p$ for the $16$-qubit instances), and moreover shows that the trained QAOA parameters give continual solution improvement as a function of $p$.
@@
The $27$ qubit instances of Figures 2 and 3 (right) were trained only up to $p=7$ because of the high classical computational cost required to train these instances up to very high rounds – the $16$ qubit instances of Figures 1 and 3 (left) were significantly easier to train because the size of the statevector simulation is significantly smaller.
@@
Figure 7: IBM quantum processor QAOA performance results for the $156$ qubit IBM quantum computers in terms of objective function (Ising model) energy on the y-axis as a function of the QAOA step size $p$ on the log scale x-axis. Each processor is sampling a single $156$-spin hardware-defined Ising model instance, whose global minimum energy is 246-246. Each plotted point (defined by a $p$ and a fixed set of transferred QAOA parameters) is showing the mean expectation value estimated using $10^{5}$ samples, except ibm_fez where $10^{6}$ samples are used. The classically trained angles are denoted in the legend by the qubit count of the problem followed by an index integer of the random instance (or all positive or negative coefficients), resulting in a total of $24$ sets of QAOA angles. Within each sub-plot, the red and blue lines each show the two sets of transferred QAOA angles, for $27$ and $16$-qubit source instances, which resulted in the best (lowest) expectation value found on that particular QPU. Although these 55 sub-plots show sampling results from the same fixed Ising model, the y-axis are not scaled to be the same for each device so as to highlight any small differences between the noisy computations.
@@ III.2 Validating Small-Scale Parameter Transfer with Classical Statevector Simulations
Figure 4 shows, as a function of $p$, what the mean energy is when using a fixed set of QAOA angles learned from one of the (random-coefficient) $16$-qubit instances on all of the other $16$ and $27$-qubit problem instances (which, this set of fixed angles was specifically not trained on). These energy curves remarkably show that despite several blips of non-improving energy as a function of $p$, there is a very consistent energy improvement up to very large $p$. Fundamentally, this shows that the limits of parameter transfer, at least for this size of problem instance (in terms of what size of $p$ can still be used effectively) is up to the saturation of an approximation ratio of $1$ on the training instance – if you are willing to accept temporary non-ideal blips of non consistent improvement as a function of $p$.
Figure 5 shows parameter transfer performance for another set of fixed trained QAOA angles, specifically trained on the $16$-qubit Ising model with only negative coefficients. Figure 6 shows the same thing but for the $16$-qubit Ising model with all positive coefficients.
Figures 4 and 5 and 6 all show that the different trained QAOA angles do transfer onto other problem instances. In particular at sufficiently large $p$, these parameters do eventually show continuing improvement in energy as a function of $p$. This general finding that despite very different random coefficient $\{+1,-1\}$ distributions, the QAOA angles still can be effectively transferred.
@@
Figure 8: IBM quantum processor QAOA performance results in terms of objective function (Ising model) energy on the y-axis as a function of the QAOA step size $p$ on the log scale x-axis, where the parameters used for each $p$ are the same angles that were learned on the smaller problem instances classically. This is a continuation of Figure 7 using the same format, but for different IBM devices. The classically trained angles are denoted by the qubit count of the problem followed by an index integer of the random instance (or all positive or negative coefficients), resulting in a total of $24$ sets of QAOA angles. The optimization problem is a minimization task, meaning that lower energies are better solutions. Each point (defined by a $p$ and a fixed set of transferred QAOA parameters) is showing the mean energy sampled from $10^{6}$ measurements, with the exceptions of ibm_torino where only $10^{5}$ were measured. The set of angles which resulted in the lowest (best) mean energy for both the $16$-qubit and the $27$-qubit instances are shown as blue and red lines respectively. The minimum (mean) energies, with continuous energy improvement as a function of $p$, are marked with $\ast$ symbols and vertical lines at that $p$ index. The optimization problem being solved by each QPU is identical for each QPU size (since the hardware two-qubit entangling gate connectivity structure is exactly the same for each QPU size); the instance is the same for all of the $127$ qubit devices, and there is a single $133$ qubit instance for ibm_torino.
@@
Figure 9: IBM quantum processor hardware error rates, aggregated from the vendor-provided calibration data for all QAOA circuits that were executed and whose results are shown in Figure 8. Top row shows the single and two qubit gate randomized benchmarking error rates (log scale x-axis), bottom row shows T1 and T2 coherence times (x-axis is in microseconds). Each plot presents the cumulative distribution function (y-axis) of the combined hardware measured error rates for each device.
@@
Having trained an ensemble of $24$ sets of QAOA “schedules”, one approach could be to compute average quantities for these learned angles, which could represent a good schedule for an average problem. However, here we utilize all $24$ sets of angles to sample the underlying heavy-hex Ising models in order to obtain a spectrum of QAOA performance – where some outlier angles perform worse, but potentially others perform better. This is seen very clearly in Figures 4, 5, and 6 where the performance differs significantly across the different fixed angles.
@@ III.3 IBM Quantum Computer Sampling Using the Transferred QAOA Angles
Figures 7 and 8 plots the finite-shot estimates of energy expectation value as a function of the (learned) increasing $p$ angles when executed on $127$, $133$, and $156$ qubit IBM QPUs. There are only two differences between the $24$ distinct connected lines in the scaling plots as a function of $p$: i) Time, meaning there is possible time dependent noise drift on the quantum hardware (which does certainly occur on these QPUs [86, 87]), and ii) Performance differences because of the difference in the trained QAOA angles. The ensemble of learned angles are all unique (although, they are typically quite similar) because they were all trained on distinct random heavy-hex spin glass instances. In other words, the underlying combinatorial optimization problem instance that is being optimized is fixed, as is the QPU, as is the structure of the circuit being defined strictly on the heavy-hex graph (only single qubit gate rotations are changed across the different compiled circuits from the different sets of angles). Therefore, the intention is for these plots to show the relative differences between the IBM QPUs across the set of $24$ distinct sets of QAOA parameters (i.e., QAOA schedules) as the QAOA circuit depth increases, up to the largest parameter-transferred $p$-depth. There is not a single source problem instance which resulted in trained QAOA angles that resulted in the best performance on NISQ hardware, instead Figures 8 and 7 show that the on-hardware performance is quite similar across the different sets of angles (this is very likely due to the error rates, and likely biased errors, on the QPUs).
@@
Figure 8 shows that there is a clear and consistent loss of signal where the mean expectation value converges to an energy of 0 for the $127$ qubit processors when $p$ is sufficiently large (around $p=30$). Remarkably however, both the $133$ and $156$ qubit processors (see Figure 7) show a non-zero signal up to the deepest QAOA circuits tested, which were up to $p=49$. This is an indication that the qubit coherence times are significantly better for the Heron processors compared to the $127$ qubit Eagle processor devices. Although these extremely high depth QAOA circuits at $p=49$ do not show an improvement of mean expectation value, these are the highest depth QAOA digital circuits executed on noisy quantum processor to date – and it is notable that there is a nonzero signal generated by some, but not all, of the IBM quantum processors. It is surprising that despite the peak performance of the $156$-qubit IBM Heron r2 and r3 devices being at $p\approx 10$, there is a non-random Hamiltonian expectation found all the way up to $p=49$. For quantum circuit size context, the largest one of these QAOA circuits (both with respect to qubit count and circuit depth) uses $156$ qubits, $17,248$ CZ two-qubit gates, $40,924$ SX single qubit gates, $34,513$ RZ single qubit gates, and $156$ qubit measurement operations at the end of the circuit. This qualitative result is made especially clear by the ibm_pittsburgh data, where the largest QAOA circuit expectation value at $p=49$ is a lower energy (better) than, or approximately equal to, the $p=1$ results.
@@
Notably, none of the NISQ computer QAOA distributions shown in Figure 8 sampled a globally optimal energy spin configuration, despite the usage of a wide range of trained parameter transfer angles, in addition to high measurement counts. Table 2 reports the lowest energies sampled on each QPU.
@@
In order to show the component level errors rates on these IBM NISQ processors used in Figure 7 and Figure 8, Figure 9 shows aggregate single and two qubit gate error rates, as well as the qubit coherence time metrics of $T1$ and $T2$ times across most of the IBM NISQ computers. $T1$ is the relaxation time and $T2$ is the dephasing time [88] – these metrics give estimates for how long the qubits remain coherent (the larger these times are, the better the qubits are). These error rates are reported by the device backends, and here we are showing the cumulative distribution function of the distribution of all error rates collected from every job that was submitted to the device (and that successfully executed) whose results are shown in Figure 8 – this means that the reported distribution technically contains duplicates because the IBM devices are only calibrated once every some set number of hours (dependent on the device), and these metrics are reported by this device calibration on the backend. The goal of showing this distribution in this way is because it shows the aggregate effective error rates encountered by the QAOA circuits seen in Figure 8, and moreover show differences in qubit coherence times on the different hardware platforms. These metrics are measured using randomized benchmarking [89, 90] techniques. Note that ibm_nazca and ibm_kyoto are not included in this dataset because the hardware level error rates could not be extracted from the Qiskit API post-circuit execution. The gate error rates shown in Figure 9 are based on recording the error rates of the gateset to which the circuits are compiled to. The error rates of the single qubit gates are recorded based on the gates sx and x, and the error rates of the two qubit gates are recorded based on the gates ecr and cz. Note that the single qubit gate rz is a virtual gate which is implemented as an abstract circuit compilation operation [91], which means that the error rate for this gate is always 0. When retrieving two qubit gate errors rates from the IBM backends, there are many instances of reported error rates of $1.0$. It seems that these reported error rates are indicating either an uncalibrated two qubit gate, or a gate which has been de-activated at the software level. Because these error rates are an erroneous backend related artifact, we have discarded all $1.0$ error rates in the data shown in Figure 9.
@@
Figure 9 shows that the $T1$ and $T2$ times of the newer generations of IBM processors, such as ibm_marrakesh, are smaller than the previous generation QPUs. However, this does not mean that the newer devices actually have worse performance – the reason is because the single and two qubit gate times were also reduced, along with the coherence times, on these Heron processors, and the relevant measure therefore becomes circuit depth that can be executed within the coherence times.
@@
One way of quantifying the size of QAOA circuits that have been executed on noisy quantum computers such that there is continual improvement at each step of p is a concept called “QAOA volume” [11, 10] which defines the largest QAOA circuit from $N$ variables and depth $p$ that results in monotonic improvement of solutions up to that depth $p$ (analogous to the quantum volume NISQ computer benchmark [92, 93, 94, 95]). The QAOA volumes reported in the hardware simulations shown in Figures 7 and 8 are summarized in Table 2. These are some of the largest QAOA volumes reported on noisy quantum computers to date, going up to a continual improvement of $p=10$ on a $156$ qubit processor.

### v2

@@ III.1 Classical QAOA Angle Learning with Statevector Simulations
Fig. 1 plots the numerical quantities of the learned QAOA angles $\beta,\gamma$ for the ensemble of $12$ different $16$-qubit heavy-hex spin glass instances, learned until $p=49$ or the mean (noiseless) energy is at least $0.01$ (energy) below the optimal ground-state energy. During training, the QAOA angles are not restricted or bounded to a certain range. Fig. 2 plots the same but for the ensemble of $27$ qubit instances. This set of $24$ QAOA angles are the parameters that are used in the entirety of this study for evaluating how well, or not, parameter transfer works up to these very high depth QAOA circuits.
Figure 4: Numerical simulation of parameter transfer of one of the $16$ qubit problem instances onto the other $11$ $16$-qubit problem instances (left) and onto the $12$ $27$-qubit problem instances (right). These simulations are all noiseless (no shot noise). y-axis is the approximation ratio for the given target problem instances, x-axis is $p$ QAOA steps. Each line corresponds to a different problem instance (which were not using in the training of this set of fixed angles), which is indexed in the legend numerically and by the all positive and negative coefficient Ising models. Red asterisk markers denote the $p$ step where the expectation value decreases in quality (increases in energy) for that given problem instance, if that occurs. 88 out the $11$ instances in the left plot have at least one non-improving $p$ step, and 99 out of $12$ of the instances in the right plot have at least one non-improving $p$ step. However, there is an overall trend for all problem instances shown in these plots that the higher $p$ angles do give consistently improving energies, even if there are some blips of non-monotonic improvements.
Our trained angles all lie in the range $0,2\pi)$, however, there are several periodicities and symmetries that enable us to further transfer the angles into mathematical equivalent angles plotted in Figs. 12:
@@
  • Lastly, some of the trained QAOA angles have very different behavior that is not seen in the other problem instances – namely, instance number 5 in Fig. 1 shows highly fluctuating angles for the largest (converged) $p$ that was found. This highly fluctuating instance could be related to the convergence to nearly an approximation ratio of $1$.
@@
More generally, these optimized QAOA schedules closely follow recent studies that have drawn connections between highly optimized QAOA trajectories, and continuous-time quantum annealing schedules, which typically result in smooth QAOA angle trajectories [85, 86, 50, 87, 88] (this connection can be understood as QAOA being a type of discretized continuous-time adiabatic quantum annealing). These highly optimized QAOA angles, shown in Figs. 1 and 2, therefore provide more evidence for this fundamental connection between continuous time adiabatic evolution and the discretized QAOA algorithm. This smooth-schedule property could be due to the underlying angle extrapolation technique used in the training, however because the ground-truth of optimal QAOA angles are not known, there is no way to know whether these angles are optimal or not and therefore we consider the JuliQAOA angles to be good heuristic QAOA control parameters. Whether the QAOA schedule characteristics we observe are due an artifact of angle extrapolation or not, their performance on large-scale problem instances shows that these types of QAOA schedules work very well, supporting similar findings from prior studies [85, 86, 50].
@@
Fig. 3 plots the mean noiseless approximation ratios for the set of trained QAOA angles shown in Figs. 1 and 2, evaluated on the problem instances on which the QAOA parameters were trained. Fig. 3 shows the rate of energy improvement as a function of $p$ (which quickly plateaus at large $p$ for the $16$-qubit instances), and moreover shows that the trained QAOA parameters give continual solution improvement as a function of $p$.
@@
The $27$ qubit instances of Figs. 2 and 3 (right) were trained only up to $p=7$ because of the high classical computational cost required to train these instances up to very high rounds – the $16$ qubit instances of Figs. 1 and 3 (left) were significantly easier to train because the size of the statevector simulation is significantly smaller.
@@
Figure 5: IBM quantum processor QAOA performance results for the $156$ qubit IBM quantum computers in terms of objective function (Ising model) energy on the y-axis as a function of the QAOA depth $p$ on the log scale x-axis. Each processor is sampling a single $156$-spin hardware-defined Ising model instance, whose global minimum energy is 246-246. Each plotted point (defined by a $p$ and a fixed set of transferred QAOA parameters) is showing the mean expectation value estimated using $10^{5}$ samples, except ibm_fez where $10^{6}$ samples are used. The classically trained angles are denoted in the legend by the qubit count of the problem followed by an index integer of the random instance (or all positive or negative coefficients), resulting in a total of $24$ sets of QAOA angles. Within each sub-plot, the red and blue lines each show the two sets of transferred QAOA angles, for $27$ and $16$-qubit source instances, which resulted in the best (lowest) expectation value found on that particular QPU. Although these 55 sub-plots show sampling results from the same fixed Ising model, the y-axis are not scaled to be the same for each device so as to highlight any small differences between the noisy computations. Note that each figure legend contains multiple lines that are either solid ($16$-qubit training instance) or dashed ($27$-qubit training instance) gray – these lines are not visually differentiable, and are intended to show the collective sampling rates for the sets of fixed QAOA angles. The minimum (mean) energies, with continuous energy improvement as a function of $p$, are marked with $\times$ symbols and vertical lines at that $p$ index.
@@ III.2 Validating Small-Scale Parameter Transfer with Classical Statevector Simulations
Fig. 4 shows, as a function of $p$, what the mean energy is when using a fixed set of QAOA angles learned from one of the (random-coefficient) $16$-qubit instances on all of the other $16$ and $27$-qubit problem instances (which, this set of fixed angles was specifically not trained on). These energy curves remarkably show that despite several blips of non-improving energy as a function of $p$, there is a very consistent energy improvement up to very large $p$. Fundamentally, this shows that the limits of parameter transfer, at least for this size of problem instance (in terms of what size of $p$ can still be used effectively) is the saturation of an approximation ratio of $1$ on the training instance – if you are willing to accept temporary non-ideal blips of non consistent improvement as a function of $p$. We argue that parameter transfer works if the performance (computed expectation value) is monotonically improving as a function of $p$; this is the same requirement we impose on the initial QAOA training. However, if this condition is not met, parameter transferred angles could still work well if the performance on average tends to improve at larger $p$ relative to smaller $p$ despite transitory decreases in performance.
@@
Figure 6: IBM quantum processor QAOA performance results in terms of objective function (Ising model) energy on the y-axis as a function of the QAOA depth $p$ on the log scale x-axis, where the parameters used for each $p$ are the same angles that were learned on the smaller problem instances classically. This is a continuation of Fig. 5 using the same format, but for different IBM devices. The classically trained angles are denoted by the qubit count of the problem followed by an index integer of the random instance (or all positive or negative coefficients), resulting in a total of $24$ sets of QAOA angles. The optimization problem is a minimization task, meaning that lower energies are better solutions. Each point (defined by a $p$ and a fixed set of transferred QAOA parameters) is showing the mean energy sampled from $10^{6}$ measurements, with the exceptions of ibm_torino where only $10^{5}$ were measured. The set of angles which resulted in the lowest (best) mean energy for both the $16$-qubit and the $27$-qubit instances are shown as blue and red lines respectively. The minimum (mean) energies, with continuous energy improvement as a function of $p$, are marked with $\times$ symbols and vertical lines at that $p$ index. The optimization problem being solved by each QPU is identical for each QPU size (since the hardware two-qubit entangling gate connectivity structure is exactly the same for each QPU size); the instance is the same for all of the $127$ qubit devices, and there is a single $133$ qubit instance for ibm_torino.
@@
Figure 7: IBM quantum processor hardware error rates, aggregated from the vendor-provided calibration data for all QAOA circuits that were executed and whose results are shown in Fig. 6 and Fig. 5. Top row shows the single and two qubit gate randomized benchmarking error rates (log scale x-axis), bottom row shows T1 and T2 coherence times (x-axis is in microseconds). Each plot presents the cumulative distribution function (y-axis) of the combined hardware measured error rates for each device.
@@
Having trained an ensemble of $24$ sets of QAOA “schedules”, one approach could be to compute average quantities for these learned angles, which could represent a good schedule for an average problem. However, here we utilize all $24$ sets of angles to sample the underlying heavy-hex Ising models in order to obtain a spectrum of QAOA performance – where some outlier angles perform worse, but potentially others perform better. This is seen very clearly in Figs. 4 where the performance of the fixed angles differs significantly across the different instances.
Appendix B reports small-instance size parameter transfer numerical simulation results in Figs. 14, and 15 where the donor instances are entirely positive and entirely negative coefficient Ising models. The angles from these uniformly weighted models show a similar story where the transferred angles work well, but in some cases and at some $p$ depths the approximation ratio improvement is not strictly monotonic. Moreover, the entirely positive-weighted coefficient Ising model training angles do not perform as well when transferred to other instances compared to the all negative coefficient model.
@@ III.3 IBM Quantum Computer Sampling Using the Transferred QAOA Angles
Figs. 5 and 6 plots the finite-shot estimates of energy expectation value as a function of the (learned) increasing $p$ angles when executed on $127$, $133$, and $156$ qubit IBM QPUs. There are only two differences between the $24$ distinct connected lines in the scaling plots as a function of $p$: i) Time, meaning there is possible time dependent noise drift on the quantum hardware (which does certainly occur on these QPUs [89, 90]), and ii) Performance differences of the transferred QAOA angles. The ensemble of learned angles are all unique (although, they are typically quite similar) because they were all trained on distinct random heavy-hex spin glass instances. The underlying combinatorial optimization problem instance that is being optimized is fixed, as is the QPU, as is the structure of the circuit being defined strictly on the heavy-hex graph (only single qubit gate rotations are changed across the different compiled circuits from the different sets of angles). Therefore, the intention is for these plots to show the relative differences between the IBM QPUs across the set of $24$ distinct sets of QAOA parameters (i.e., QAOA schedules) as the QAOA circuit depth increases, up to the largest parameter-transferred $p$-depth. There is not a single source problem instance which resulted in trained QAOA angles that resulted in the best performance on NISQ hardware, instead Figs. 6 and 5 show that the on-hardware performance is quite similar across the different sets of angles (this is very likely due to the error rates, and likely biased errors, on the QPUs).
@@
Fig. 6 shows that there is a clear and consistent loss of signal where the mean expectation value converges to an energy of 0 for the $127$ qubit processors when $p$ is sufficiently large (around $p=30$). Remarkably however, both the $133$ and $156$ qubit processors (see Fig. 5) show a non-zero signal up to the deepest QAOA circuits tested, which were up to $p=49$. This is an indication that the qubit coherence times are significantly better for the Heron processors compared to the $127$ qubit Eagle processor devices. Although these extremely high depth QAOA circuits at $p=49$ do not show a continual improvement of mean expectation value (because of noise), these are the highest depth QAOA digital circuits executed on noisy quantum processor to date – and it is notable that there is a nonzero signal generated by some, but not all, of the IBM quantum processors. It is surprising that despite the peak performance of the $156$-qubit IBM Heron r2 and r3 devices being at $p\approx 10$, there is a non-random Hamiltonian expectation found all the way up to $p=49$. For quantum circuit size context, the largest one of these QAOA circuits (both with respect to qubit count and circuit depth) uses $156$ qubits, $17,248$ CZ two-qubit gates, $40,924$ SX single qubit gates, $34,513$ RZ single qubit gates, and $156$ qubit measurement operations at the end of the circuit. This qualitative result is made especially clear by the ibm_pittsburgh data, where the largest QAOA circuit expectation value at $p=49$ is a lower energy (better) than, or approximately equal to, the $p=1$ results.
@@
Notably, none of the NISQ computer QAOA distributions shown in Fig. 6 sampled a globally optimal energy spin configuration, despite the usage of a wide range of trained parameter transfer angles, in addition to high measurement counts. Table 2 reports the lowest energies sampled on each QPU.
@@
In order to show the component level errors rates on these IBM NISQ processors used in Fig. 5 and Fig. 6, Fig. 7 shows aggregate single and two qubit gate error rates, as well as the qubit coherence time metrics of $T1$ and $T2$ times across most of the IBM NISQ computers. $T1$ is the relaxation time and $T2$ is the dephasing time [91] – these metrics give estimates for how long the qubits remain coherent (the larger these times are, the better the qubits are). These error rates are reported by the device backends, and here we are showing the cumulative distribution function of the distribution of all error rates collected from every job that was submitted to the device (and that successfully executed) whose results are shown in Fig. 6 – this means that the reported distribution technically contains duplicates because the IBM devices are only calibrated once every some set number of hours (dependent on the device), and these metrics are reported by this device calibration on the backend. The goal of showing this distribution in this way is because it shows the aggregate effective error rates encountered by the QAOA circuits seen in Fig. 6, and moreover show differences in qubit coherence times on the different hardware platforms. These metrics are measured using randomized benchmarking [92, 93] techniques. Note that ibm_nazca and ibm_kyoto are not included in this dataset because the hardware level error rates could not be extracted from the Qiskit API post-circuit execution. The gate error rates shown in Fig. 7 are based on recording the error rates of the gateset to which the circuits are compiled to. The error rates of the single qubit gates are recorded based on the gates sx and x, and the error rates of the two qubit gates are recorded based on the gates ecr and cz. Note that the single qubit gate rz is a virtual gate which is implemented as an abstract circuit compilation operation [94], which means that the error rate for this gate is always 0. When retrieving two qubit gate errors rates from the IBM backends, there are many instances of reported error rates of $1.0$. It seems that these reported error rates are indicating either an uncalibrated two qubit gate, or a gate which has been de-activated at the software level. Because these error rates are an erroneous backend related artifact, we have discarded all $1.0$ error rates in the data shown in Fig. 7.
@@
Fig. 7 shows that the $T1$ and $T2$ times of the newer generations of IBM processors, such as ibm_marrakesh, are smaller than the previous generation QPUs. However, this does not mean that the newer devices actually have worse performance – the reason is because the single and two qubit gate times were also reduced, along with the coherence times, on these Heron processors, and the relevant measure therefore becomes circuit depth that can be executed within the coherence times.
@@
One way of quantifying the size of QAOA circuits that have been executed on noisy quantum computers such that there is continual improvement at each step of p is a concept called “QAOA volume” [11, 10] which defines the largest QAOA circuit from $N$ variables and depth $p$ that results in monotonic improvement of solutions up to that depth $p$ (analogous to the quantum volume NISQ computer benchmark [95, 96, 97, 98]). The QAOA volumes reported in the hardware simulations shown in Figs. 5 and 6 are summarized in Table 2. These are some of the largest QAOA volumes reported on noisy quantum computers to date, going up to a continual improvement of $p=10$ on a $156$ qubit processor. QAOA volume here shows a benchmark of current IBM NISQ computer capability.